# Combining complex conjugation, time-reversal, and spin symmetry projection of coupled cluster wave functions

### v1

@@ Abstract
Complex conjugation symmetry breaking and restoration generate two non-orthogonal configurations at the Hartree-Fock level that can capture static correlation naturally. In conjunction with broken spin-symmetry coupled cluster theory, the symmetry-projected wave function shows good agreement with full configuration interaction in beryllium hydride insertion, lithium fluoride dissociation, and symmetric stretching of tetrahedral H4. By adding spin flip projection, we can also recover time reversal symmetry in the same coupled cluster framework.
@@ I Introduction
In this work, we review the complex conjugation operator and implement the complex conjugation projection of Hartree-Fock and coupled cluster theory. In addition, we add spin flip projection to recover time-reversal symmetry. We test these methods on small molecules with complex restricted Hartree-Fock, unrestricted Hartree-Fock (UHF), and generalized Hartree-Fock, as well as with coupled cluster with single and double excitations (CCSD) based on all of these mean-field references. We aim to show that complex conjugation projection can be a useful tool for multi-reference problems. We also show a few pitfalls which might arise, since frequently projecting one symmetry also projects or at least approximately projects others, and we show how we can circumvent these difficulties.
@@ II.2 The Time Reversal and Spin Flip Operators
We introduce the spin flip operator $F=e^{i\pi S_{y}}$ and time reversal operator $\mathcal{T}=FK$. Together with the identity operator and complex conjugation operator, the operator product is closed:
@@ III.1 Complex Conjugation Projected Hartree-Fock
It is easy to show that that
@@
In other words, given a complex determinant $|\tilde{\Phi}\rangle$ which is projected by $P_{K}=1+\mathrm{e}^{\mathrm{i}\,\theta}K$ so that the projected state $P_{K}|\tilde{\Phi}\rangle$ is a $K$ eigenstate with eigenvalue $\mathrm{e}^{-\mathrm{i}\,\theta}$, we can construct an associated complex determinant $|0\rangle$ such that the projector is simply $1+K$ and the corresponding $K$ eigenvalue of $P_{K}|0\rangle$ is just 1.
@@ III.2 Connection with Other Multireference Methods
Before we discuss complex conjugation projected coupled cluster theory, let us take a moment to consider the structure of the KHF state. To simplify the presentation, we assume that we are projecting a complex RHF determinant that has only one (complex) occupied orbital $|\chi\rangle$. We can write $\chi$ in terms of two two real molecular orbitals $\phi_{p}$ and $\phi_{q}$ as
@@
After projection, only the real part survives, and it takes the form of a perfect pairing generalized valence bond state, except that $\phi_{p}$ and $\phi_{q}$ need not be individually orthonormal (but can be).
@@ III.3 Complex Conjugation Projected Coupled Cluster
After minimizing the PHF energy to get the KHF broken symmetry determinant $|\tilde{\Phi}\rangle=e^{Z}|\Phi\rangle$, we can use the procedure of Eqn. 13 to reduce the projection operator to simply $1+K$. We discuss the reason for this gauge-fixing step below. We can then add coupled cluster theory to recover the dynamic correlation. We mimic the traditional coupled cluster equations to write down projected coupled cluster equations
@@
where a * on an operator means taking the complex conjugate of its coefficients. For example, $T^{\star}$ has the usual form of the cluster operator but uses the complex conjugate of the cluster amplitudes.
@@ IV Results
Figure 1: Top panel: Energy of various HF and CCSD in LiF dissociation. Bottom panel: Energy error in LiF dissociation.
@@ IV.1 LiF Dissocation
Now we turn to the CCSD results. In this case, we use the RHF reference for smaller bond lengths and the UHF reference for larger bond lengths. The fact that the reference changes explains the discontinuity observed in the CCSD curve in the bottom panel of Fig. 1. For shorter bond lengths, KCCSD is much more accurate than CCSD, but the energy error increases significantly once spin symmetry is broken, and unrestricted CCSD and KCCSD have similar errors overall. Importantly, however, where UCCSD is not even continuous, KUCCSD yields a continuous albeit not smooth dissociation profile.
@@
Tetrahedral H4 is a small molecule with spin frustration. Due to its unique geometry, one can find a real RHF, a real UHF, a real GHF, and a complex GHF solution.[7] All mean-field solutions other than real RHF break time-reversal symmetry, so, in this case, we restore both time reversal and complex conjugation (by, as noted, projectively restoring complex conjugation and spin-flip symmetries). In the dissociation, there is a singlet-triplet crossing at short bond length. Thus, we use $1+F$ for singlet (denoted as S0) and $1-F$ for triplet (denoted as S1) projection respectively.
Figure 3 shows that at the mean-field level, the system is not bonded. Complex conjugation projected GHF (KGHF) also has no bond, though complex conjugation and singlet projected GHF (S0 KGHF) has a very small minimum at an H-H distance of about 1.9 Bohr. Also, we note that the crossing of singlet and triplet KGHF occurs for hydrogen-hydrogen bond lengths about 1.2 Bohr, at a slightly smaller bond length than the FCI singlet and triplet solutions cross.
As a system with only four electrons, the complex GHF-based CCSD (cGCCSD) is not too far from FCI, but the wave function switched from triplet to singlet at 1.25 Bohr, where the FCI solutions cross. That is, the cGCCSD tracks the lower energy FCI solution rather than remaining on a single state. The same behavior has been seen elsewhere in the GCCSD solution for O2 dissociation.[24] Generally, KGCCSD is much closer to FCI than is GCCSD, but it also tracks the lower energy FCI solution. Including time reversal projection not only yields improved energies but also allows us to partially control the spin state onto which we project. Figure 4 shows the maximum errors are about 1 and 3 milliHartree for singlet and triplet respectively. However, at sufficiently short bond length we are unable to fully control spin symmetry just by complex conjugation and time reversal projection since, as we noted, this amounts to half-spin projection. Evidently, full spin projection is needed. We note, though, that projecting both complex conjugation and time reversal amounts to a PHF which is a linear combination of 4 determinants, while fully spin-projected GHF requires integrating over three continuous Euler angles, which would generally be expected to require a much larger number of states in the PHF.
@@ IV.3 BeH2
At the FCI level, meanwhile, there are three low-energy singlets. Two are of A1 symmetry and reflect the two RHF solutions we have just described, and the third is of B2 symmetry and is essentially the open-shell singlet arising from occupied the frontier $B_{2}$ and $A_{1}$ orbitals once each. Where the RHF states cross near $x=1.5$, the FCI instead has an avoided crossing, and this avoided crossing implies strong correlation. Complicating things somewhat is that where the exact ground state is of A1 symmetry across most of the insertion pathway, the B2 open-shell singlet is the ground state near $x=1.5$.
The middle panel of Figure 5 shows the energy of various PHF states in comparison to the FCI benchmark. Comparing the top and middle panels reveals that, though complex conjugation projected RHF (KRHF) is far from exact, it is still notably more accurate than is RHF. Unfortunately, KRHF has a shoulder at $x\approx 1.35$ and another at $x\approx 1.65$. This is due to the presence of the FCI B2 solution. The complex RHF has, as its highest occupied molecular orbital, a linear combination of a1 and $b_{2}$ character so that complex RHF has broken point group symmetry. Complex conjugation projection tends also to restore point group symmetry, and the ground state in the region of the shoulder is of B2 symmetry, not of A1 symmetry as we see elsewhere. The bottom panel of Fig. 5 makes this point clearly. We see that there are three distinct KRHF solutions. For $x\lesssim 1.35$ and $x\gtrsim 1.63$, the KRHF is of A1 symmetry corresponding to the two distinct A1 RHF solutions. For $x$ between these two values, KRHF instead gives a B2 state. We can follow these three distinct KRHF solutions beyond the point at which they are the KRHF ground state, but at some point we are unable to locate them.
@@ V.1 Discussion
On the other hand, as discussed elsewhere,[12, 13] projected CC has a problem with what we call gauge invariances. These are numerical invariances of the algebraic equations that leave observables unchanged. Although gauge invariances do not affect the final result, they still make the PCCSD result not unique when using approximation methods. In this work, we choose the projection operator to be $1+K$ to fix the gauge in complex conjugation, but whether it is the best choice still needs to be determined. When introducing spin-flip symmetry, we have an extra spin gauge mode, and this mode causes trouble in the convergence of the PCC equations. The spin-flip projection only contains two operators, but we need to use large damping factors to slow down changes in CC amplitudes. Otherwise, it oscillates dramatically.
V.2 Conclusion
@@
VI Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
@@ Appendix A Gradient for projected HF
In this appendix, we provide a detailed discussion of the optimization of symmetry projected HF states based on a Thouless parametrization. In Eqn 11, we treat $z$ and $z^{*}$ as independent variables, the gradient is
$G_{ai}=\frac{\partial E}{\partial z^{*}_{ai}}=\frac{\langle\Phi|c_{i}^{\dagger}c_{a}\,e^{Z^{\dagger}}\,\overleftarrow{P_{K}}\,P^{\dagger}_{F}\,(H-E)\,P_{F}\,P_{K}\,e^{Z}|\Phi\rangle}{\langle\Phi|e^{Z^{\dagger}}\,\overleftarrow{P_{K}}\,P^{\dagger}_{F}\,P_{F}\,P_{K}\,e^{Z}|\Phi\rangle}$ (23)
Since we choose the reference state to be real, complex conjugation operator will only change the parameter $z,z^{*}$ and leave the reference unchanged.
The gradient is made of two parts, the kernel for the overlap-like part is
@@
and for the Hamiltonian-like part is
$\frac{\langle\Phi|c_{i}^{\dagger}c_{a}\,e^{Z^{\dagger}}\,H\,R\,e^{Z}|\Phi\rangle}{\langle\Phi|e^{Z^{\dagger}}\,R\,e^{Z}|\Phi\rangle}=h\,\rho_{ai}+[(1-\rho)\,(h+\Gamma)\,\rho]_{ai}$ (25)
Here $R$ is the operator in the projection operators, and $h$ and $\Gamma$ are defined as following
$h=\frac{\langle\Phi|e^{Z^{\dagger}}\,H\,R\,e^{Z}|\Phi\rangle}{\langle\Phi|e^{Z^{\dagger}}\,R\,e^{Z}|\Phi\rangle}$ (26)
$\Gamma_{ik}=\sum_{jl}V_{ijkl}\,\rho_{lj}$ (27)

### v2

# Combining complex conjugation, time-reversal, and spin-flip symmetry projection of coupled cluster wave functions

@@ Abstract
Complex conjugation symmetry breaking and restoration generates two non-orthogonal configurations at the Hartree-Fock level that can capture static correlation naturally. In conjunction with broken spin-symmetry coupled cluster theory, the symmetry-projected wave function shows good agreement with full configuration interaction in beryllium hydride insertion, lithium fluoride dissociation, and symmetric stretching of tetrahedral H4. By adding spin flip projection, we can also recover time reversal symmetry in the same coupled cluster framework. We also show results including point group symmetry projection.
@@ I Introduction
In this work, we review the complex conjugation operator and implement the complex conjugation projection of Hartree-Fock and coupled cluster theory. In addition, we add spin flip projection to recover time-reversal symmetry. We test these methods on small molecules with complex restricted Hartree-Fock, unrestricted Hartree-Fock (UHF), and generalized Hartree-Fock, as well as with coupled cluster with single and double excitations (CCSD) based on all of these mean-field references. We aim to show that complex conjugation projection can be a useful tool for multi-reference problems. We also show a few pitfalls which might arise, since frequently projecting one symmetry also projects or at least approximately projects others, and we show how we can circumvent these difficulties by adding extra projectors to cleanly disambiguate states. In particular, projecting complex conjugation can also restore point group symmetry, and to have full control over the final states we can include point group symmetry projection along with all the others.
@@ II.2 The Time Reversal and Spin Flip Operators
We now turn to the spin flip operator $F=e^{i\pi S_{y}}$ and time reversal operator $\mathcal{T}=FK$, where in its action on electrons we may write
$S_{y}=\frac{1}{2\,\mathrm{i}}\,\sum_{p}\left(c_{p,\uparrow}^{\dagger}\,c_{p,\downarrow}-h.c.\right).$ (6)
Pragmatically, $F$ converts a determinant with orbital coefficients $\left(\begin{smallmatrix}\bm{C}_{\uparrow}\\ \bm{C}_{\downarrow}\end{smallmatrix}\right)$ into one with orbital coefficients $\left(\begin{smallmatrix}\bm{C}_{\downarrow}\\ -\bm{C}_{\uparrow}\end{smallmatrix}\right)$.
Together with the identity operator and complex conjugation operator, the operator product is closed:
@@ III.1 Complex Conjugation Projected Hartree-Fock
It is easy to show that
@@
In other words, given a complex determinant $|\tilde{\Phi}\rangle$ which is projected by $P_{K}=1+\mathrm{e}^{\mathrm{i}\,\theta}K$ so that the projected state $P_{K}|\tilde{\Phi}\rangle$ is a $K$ eigenstate with eigenvalue $\mathrm{e}^{-\mathrm{i}\,\theta}$, we can construct an associated complex determinant $|0\rangle$ such that the projector is simply $1+K$ and the corresponding $K$ eigenvalue of $P_{K}|0\rangle$ is just 1. We will take advantage of this freedom when projecting the coupled cluster equations, where it slightly simplifies the formalism.
@@ III.2 Connection with Other Multireference Methods
Before we discuss complex conjugation projected coupled cluster theory, let us take a moment to consider the structure of the KHF state. To simplify the presentation, we assume that we are projecting a complex RHF determinant that has only one (complex) occupied orbital $|\chi\rangle$. We can write $|\chi\rangle$ in terms of two real molecular orbitals $|\phi_{p}\rangle$ and $|\phi_{q}\rangle$ as
@@
After projection, only the real part survives, and it takes the form of a perfect pairing generalized valence bond state,[22, 23, 24, 25, 26, 27] except that $|\phi_{p}\rangle$ and $|\phi_{q}\rangle$ need not be individually orthonormal (but can be). Once we have more orbitals involved, the picture is more complicated, and in general KHF has less flexibility than a perfect pairing valence bond state for multiple electron pairs.
@@ III.3 Complex Conjugation Projected Coupled Cluster
After minimizing the PHF energy to get the KHF broken symmetry determinant $|\tilde{\Phi}\rangle=e^{Z}|\Phi\rangle$, we can use the procedure of Eqn. 14 to reduce the projection operator to simply $1+K$. We discuss the reason for this gauge-fixing step below. We can then add coupled cluster theory to recover the dynamic correlation. We mimic the traditional coupled cluster equations to write down projected coupled cluster equations
@@
where a on an operator means taking the complex conjugate of its coefficients. For example, $T^{\star}$ has the usual form of the cluster operator but uses the complex conjugate of the cluster amplitudes.
@@
Finally, we must say a few words about the projector $P_{K}$. In general, the phase factor $\mathrm{e}^{\mathrm{i}\,\theta}$ for the projected Hartree-Fock and projected coupled cluster wave functions will differ. We can solve the KCC equations while adjusting this phase or while leaving it the same as it is in KHF. When the KCC amplitude equations are converged, we find that these two choices are equivalent.
@@ IV Results
Figure 1: Energy of various HF and CCSD in LiF dissociation.
Figure 3: Top panel: two KRHF solutions for the dissociation of LiF. Bottom panel: Energy error in LiF dissociation with CCSD methods.
@@ IV.1 LiF Dissocation
One might wonder whether KRHF, as a two-determinant wave function, is able to correctly dissociate LiF. This appears not to be the case (see top panel of Fig. 3). The lowest energy KRHF solution at equilibrium dissociates to ionic fragments instead of neutral ones. There is a KRHF curve that dissociates to neutral fragments, but it is above the ionic solution except for large bond lengths and we are unable to follow it inward to less than about 4.7 Å. On the other hand, KUHF in this case appears to be size consistent (data not shown).
Now we turn to the CCSD results. In this case, we use the RHF reference for smaller bond lengths and the UHF reference for larger bond lengths. The fact that the reference changes explains the discontinuity observed in the CCSD curve in the bottom panel of Fig. 3. For shorter bond lengths, KCCSD is much more accurate than CCSD, but the energy error increases significantly once spin symmetry is broken, and unrestricted CCSD and KCCSD have similar errors overall. Importantly, however, where UCCSD is not even continuous, KUCCSD yields a continuous albeit not differentiable dissociation profile (see inset in the bottom panel of Fig. 3).
@@
Tetrahedral H4 is a small molecule with spin frustration. Due to its unique geometry, one can find a real RHF, a real UHF, a real GHF, and a complex GHF solution.[7] The real RHF solution is an $S^{2}$ eigenstate with doubly-occupied spatial orbitals, while the real UHF has different orbitals for different spins, so respects $S_{z}$ symmetry but not $S^{2}$. The real GHF solution breaks both $S^{2}$ and $S_{z}$ symmetries but yields coplanar magnetism, while the complex GHF solution yields a noncoplanar magnetic structure. Except for the real RHF solution, all of the mean-field solutions we found break time-reversal symmetry, so, in this case, we restore both time reversal and complex conjugation (by, as noted, projectively restoring complex conjugation and spin-flip symmetries). We note that paired mean-field solutions which break complex conjugation symmetry but preserve time reversal are possible in principle,[2, 3] but we have not found them here.
In the dissociation, there is a singlet-triplet crossing at short bond length. Thus, we use $1+F$ for singlet (denoted as $S_{0}$ because the $S$ quantum number for a singlet is 0) and $1-F$ for triplet (denoted as $S_{1}$ since the $S$ quantum number for a triplet is 1) projection respectively.
@@
Figure 4 shows that at the mean-field level, the system is not bonded. Complex conjugation projected GHF (KGHF) also has no bond, though complex conjugation and singlet time-reversal projected GHF (KTGHF $S_{0}$)) has a very small minimum at an H-H distance of about 1.9 Bohr. Also, we note that the crossing of singlet and triplet KTGHF (the latter denoted as KTGHF $S_{1}$) occurs for hydrogen-hydrogen bond lengths about 1.2 Bohr, at a slightly smaller bond length than the FCI singlet and triplet solutions cross.
As a system with only four electrons, the complex GHF-based CCSD (cGCCSD) is not too far from FCI, but the wave function switches from triplet to singlet at 1.25 Bohr, where the FCI solutions cross. That is, the cGCCSD tracks the lower energy FCI solution rather than remaining on a single state. Similar behavior has been seen elsewhere in the GCCSD solution for O2 dissociation.[30] Generally, KGCCSD is much closer to FCI than is GCCSD, but it also tracks the lower energy FCI solution. Including time reversal projection not only yields improved energies but also allows us to partially control the spin state onto which we project. Figure 5 shows the maximum errors are about 1 and 3 mEh for singlet and triplet respectively. However, at sufficiently short bond length we are unable to fully control spin symmetry just by complex conjugation and time reversal projection since, as we noted, this amounts to only half-spin projection. Evidently, full spin projection is needed. We note, though, that projecting both complex conjugation and time reversal amounts to a PHF which is a linear combination of 4 determinants, while fully spin-projected GHF requires integrating over three continuous Euler angles, and would generally be expected to require a much larger number of states in the PHF.
@@ IV.3 BeH2
At the FCI level, meanwhile, there are three low-energy singlets. Two are of A1 symmetry and reflect the two RHF solutions we have just described, and the third is of B2 symmetry and is essentially the open-shell singlet arising from occupying the frontier B2 and A1 orbitals once each. Where the RHF states cross near $x=1.5$, the FCI instead has an avoided crossing, and this avoided crossing implies strong correlation. Complicating things somewhat is that where the exact ground state is of A1 symmetry across most of the insertion pathway, the B2 open-shell singlet is the ground state near $x=1.5$.
The middle panel of Figure 6 shows the energy of various PHF states in comparison to the FCI benchmark. Comparing the top and middle panels reveals that, though complex conjugation projected RHF (KRHF) is far from exact, it is still notably more accurate than is RHF. Unfortunately, KRHF has a shoulder at $x\approx 1.35$ and another at $x\approx 1.65$. This is due to the presence of the FCI B2 solution. The complex RHF has, as its highest occupied molecular orbital, a linear combination of a1 and b2 character so that complex RHF has broken point group symmetry. Complex conjugation projection tends also to restore point group symmetry, and the ground state in the region of the shoulder is of B2 symmetry, not of A1 symmetry as we see elsewhere. The bottom panel of Fig. 6 makes this point clearly. We see that there are three distinct KRHF solutions. For $x\lesssim 1.35$ and $x\gtrsim 1.63$, the KRHF is of A1 symmetry corresponding to the two distinct A1 RHF solutions. For $x$ between these two values, KRHF instead gives a B2 state. We can follow these three distinct KRHF solutions beyond the point at which they are the KRHF ground state, but at some point we are unable to locate them.
@@ V.1 Discussion
On the other hand, as discussed elsewhere,[12, 33, 34, 13] projected CC has a problem with what we call gauge invariances. These are numerical invariances of the algebraic equations that leave observables unchanged. Although gauge invariances do not affect the final result, they still make the PCCSD result not unique when the projection is approximated (by, for example, truncating the operator $W$ to exclude triple excitations and higher). Partial solutions have been proposed,[12, 33, 34] but no fully satisfactory means of eliminating them has been found.
In this work, we choose the projection operator to be $1+K$ to fix the gauge in complex conjugation, but whether it is the best choice still needs to be determined. When introducing spin-flip symmetry, we have an extra spin gauge mode, and this mode causes trouble in the convergence of the PCC equations. The spin-flip projection only contains two operators, but we need to use large damping factors to slow down changes in CC amplitudes which otherwise oscillate dramatically.
V.2 Conclusions
@@ Appendix A Gradient for Projected HF
In this appendix, we detail the optimization of symmetry-projected HF states based on a Thouless parametrization. We treat the Thouless parameters $z$ and $z^{\star}$ in Eqn 12 as independent variables, and the gradient becomes
$G_{ai}=\frac{\partial E}{\partial z^{*}_{ai}}=\frac{\langle\Phi|c_{i}^{\dagger}c_{a}\,e^{Z^{\dagger}}\,\overleftarrow{P_{K}}\,P^{\dagger}_{F}\,(H-E)\,P_{F}\,P_{K}\,e^{Z}|\Phi\rangle}{\langle\Phi|e^{Z^{\dagger}}\,\overleftarrow{P_{K}}\,P^{\dagger}_{F}\,P_{F}\,P_{K}\,e^{Z}|\Phi\rangle}.$ (24)
Since we choose the reference state to be real, the complex conjugation operator will only change the parameters $z,z^{*}$ and leave the reference unchanged.
The energy gradient is made of two parts. The kernel for the overlap-like part is
@@
and that for the Hamiltonian-like part is
$\frac{\langle\Phi|c_{i}^{\dagger}c_{a}\,e^{Z^{\dagger}}\,H\,R\,e^{Z}|\Phi\rangle}{\langle\Phi|e^{Z^{\dagger}}\,R\,e^{Z}|\Phi\rangle}=h\,\rho_{ai}+[(1-\rho)\,(h+\Gamma)\,\rho]_{ai}.$ (26)
Here $R$ is the operator in the projection operators, and $h$ and $\Gamma$ are defined as
$h=\frac{\langle\Phi|e^{Z^{\dagger}}\,H\,R\,e^{Z}|\Phi\rangle}{\langle\Phi|e^{Z^{\dagger}}\,R\,e^{Z}|\Phi\rangle},$ (27)
$\Gamma_{ik}=\sum_{jl}V_{ijkl}\,\rho_{lj}.$ (28)